May 6, 2015 Calibration
GlueX Calibration Meeting
Wednesday, May 6, 2015
11:00 am, EDT
JLab: CEBAF Center, F326
Connection Using Bluejeans
- To join via Polycom room system go to the IP Address: 188.8.131.52 (bjn.vc) and enter the meeting ID: 630804895.
- To join via a Web Browser, go to the page  https://bluejeans.com/630804895.
- To join via phone, use one of the following numbers and the Conference ID: 630804895
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
- More information on connecting to bluejeans is available.
- Collaboration Meeting
- Commissioning Data
- Data quality
- Mode 7 v. Mode 8
- Cosmic data
- Subdetector Reports
- Simulations (Sean/Mark I.)
- Calibration Processing
Attending: Sean (NU); Curtis, Naomi (CMU); Noemi (Regina); Simon, Justin, Paul, Nathan, Alex, Eric, Mark I., Mike S., Will M., Adesh, Mark D., Luke
There will be an offline monitoring run for the Fall 2014 and Spring 2015 data sets tomorrow (5/7) at 5 PM EST [time changed in press]. Please have all of your updates in by then.
Justin reviewed his latest results on looking for the effects of linear polarization in the spring data, which look very good.
We reviewed several aspects of the recent data:
- Towards the end of production run 2931, a problem was discovered with the data being reported by the fADC125s. It affected the files for this run starting with #93. The problem is still being studied and it is not clear how much other data is affected.
- The data taken in the spring was a mix of readout modes 7 and 8 (without and with waveform information), with the intent of verifying that they give consistent results. Simon reported that he didn't see any gross differences between the two, though he hadn't looked at the results in detail. He did notice that in at least one run, ~10% of events had FCAL hits without pedestal information. Mike found that the error handling for fADC125 mode 7 data in the offline code required some improvements, and checked these in.
- We discussed plans for taking cosmic data. The consensus was that using the BCAL trigger was fine, and as much data as possible should be taken. This data would be useful for the BCAL, CDC, FDC (to some extent), and Start Counter. It is not clear if any other detectors would find cosmic ray data useful.
Adesh presented some preliminary results on FCAL pi0's from Run 2931 with different cluster energy thresholds. Note that these distributions differ from the monitoring plots only by excluding clusters near the center, which primarily removes a low-energy enhancement near threshold. Some concerns about the fits shown were raised, since the steeply-varying background could affect the results for fitting the peak, which gives a pi0 mass ~10-20 MeV lower than expected. Adesh will look at these and update his results for the full spring run.
He also plans to upload some updated FCAL gains, and will write a logbook entry describing what was done.
Will presented some BCAL pi0 plots from the spring run with different energy thresholds. Clear pi0 peaks were seen with consistent masses. A hint of an eta peak is seen as well, work to reduce the backgrounds in that region is on-going.
The pi0 yield is about 75% of that from the Fall run. It was suggested to look at the straight track for pi0's as well.
Mike is looking at the per-straw efficiency, since there was not enough low-noise data taken to determine the full tracking efficiency. He and Simon are also looking at the gas mix. Several methods indicate that the CO2 percentage is ~48%, and the drift tables were updated accordingly.
Lubomir is working on updating the FDC alignment with the straight track data from the last run. Simon has added preliminary alignment constants to the CCDB.
There was a long discussion about calibrations at yesterday's TOF meeting. There are parallel efforts ongoing at FSU and JLab to calibrate the TDC times. Beni has some preliminary The improved determination of the CAEN TDC bin size also improving the calibration procedures. Mike and Simon are coordinating the alignment of the TOF times with those of other detectors.
Nathan is busy with the analysis of the PS trigger data from the spring for trigger and PS calibrations. Preliminary time offsets look similar to those from the previous run. Some PS triggers were taken in combination with other triggers - Nathan has not looked at this data yet.
The PS magnet current was also varied to get the coherent peak to show up in the PS online display. The offline PS energy binning needs to be updated to reflect this.
Alex has been busy analyzing the data from his bias voltage studies. By increasing the bias voltage, he has brought the number of fibers that meet specifications (200 pixels per pulse) from 7 to 107 out of 500. These improved bias voltages were used starting at run 3180.
Anecdotally, the online scalar display shows more even results, and was used to see the coherent peak.
Paul continues to work on the RF signal calibration.
Kei reported that some first data was taken with the polarimeter, and they are in the process of checking out their detector.
Mark I. and Sean have set up the conditions and framework to run simulations corresponding to the spring run. Mark asked about the conditions required (magnet current/radiator), and it was decided to run:
- 1300 A - diamond and amorphous radiators
- 1200 A - amorphous radiator
- 800 A - diamond radiator
- No field - amorphous radiator
Since calibration work over the summer will need to be done using 2014 and 2015 data, there have also been requests to run some more simulations for the spring data. Some software work is needed for these. Sean asked for any special requests for these, and there were none. At the very least, there have been enough changes to the reconstruction code that this will be re-run for all of the simulated data from the fall 2014 run.
We briefly discussed the possibility of running calibration jobs, either in parallel with the monitoring jobs, or as a separate train. This could be for production or development of calibrations, and would give some improvement in efficiency of running over the data. There were no complaints, so Sean will put together a proposal for the next meeting.